Operation Summary for July 2, 2010 - July 7, 2010 |
The Integrated Luminosity for the period from 6/28/10 to 7/5/10 was 59.03 pb-1. Friday July 2 The day shift began with Operations monitoring the Tevatron (TeV) store 7945 (19.46E30) and the Recycler stash (361.8E10), with the Antiproton Source (Pbar) stacking (15.2E10), with the Linac, Booster and the Main Injector (MI) operating normally, with the NuMI, MiniBooNE, and Meson FTTB T-977 experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams. During the day shift, Operations terminated store 7945 at 8:30 AM, put the TeV into shot setup, and established store 7946 at 9:30 AM with an average initial luminosity (AIL) of 321.7E30. CHL personnel investigated problems with the Pbar dry engine and discovered that the engine needed a complete overhaul. They began their repairs at 9:30 AM; this resulted in degraded stacking for a while. During the evening shift, Pbar resumed normal stacking around 8 PM. There was minimal downtime this shift. Operations disabled the T-977 beam at 6 PM. Saturday July 3 The midnight shift began with Operations monitoring store 7946 (71.39E30) and the stash (254.7E10), with Pbar stacking (14.9E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations terminated store 7946 at 5:08 AM, put the TeV into shot setup, and established store 7947 at 6:28 AM with an AIL of 329.5E30. Operations established beam to T-977 at 4 AM. During the day shift, the I- Source started to degrade; operators tuned and brought it back to its normal values. During the evening shift, Operations disabled the T-977 beam shortly after 6 PM. A Booster East Gallery rooftop AC unit tripped off and was reset by the Duty Electrician. The TeV sector E0 electronics room temperature came into alarm; FESS personnel installed a portable AC unit to cool the room. Operations reported some downtime due to various Linac and Booster RF trips. Operations terminated store 7947 at 8:11 PM, put the TeV into shot setup, and established store 7948 at 9:19 PM with an AIL of 328.5E30. Sunday July 4 The midnight shift began with Operations monitoring store 7948 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations established beam to T-977 shortly after 4 AM. During the day shift, Operations terminated store 7948 at 11:47 AM, put the TeV into shot setup, and established store 7949 at 1:21 PM with an AIL of 316E30. Operators tuned the I- Source early in the shift. The MI-10 equipment room temperature came into alarm and FESS personnel were able to recover the AC unit. During the evening shift, Operations disabled the T-977 beam at 6 PM. Monday July 5 The midnight shift began with Operations monitoring store 7949 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam. During the midnight shift, scheduled beam power reductions, in preparation for the summer shutdown, began. Operations terminated store 7949 at 4:42 AM, put the TeV into shot setup, and established store 7950 at 5:53 AM with an AIL of 333.6E30. Operations established beam to T-977 once the store was spinning. The I- Source output started to drop off at the end of the shift. During the day shift, operators worked to recover the I- Source; this effort lasted throughout the shift. Operations kept the required beam intensities during the tuning. During the evening shift, Operations disabled T-977 shortly after 6 PM. The temperature of Pelletron tank 4 came into alarm. FESS personnel accessed MI-31 and found a failed AC unit due to a condenser coil clogged with vegetative debris. They cleaned all AC unit condenser coils for MI-31 and the Pelletron was operational by 8:50 PM. Pbar stacked, but the antiprotons were held in the Accumulator until FESS personnel completed the Pelletron AC work. Operations terminated store 7950 at 10:29 PM, put the TeV into shot setup, and established store 7951 at 11:59 PM with an AIL of 283.75E30. Tuesday July 6 The midnight shift began with Operations monitoring store 7951 (247.58E30) and the stash (50.4E10), with Pbar stacking (9.2E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, a MI RF station (MIRF1) tripped its filament supply breaker and would not reset; operators bypassed the station. Operations reported some downtime for NuMI when the MI-65 chillers tripped off and took down the dehumidifiers. Operations established beam to T-977 at 4:37 AM. During the day shift, Operations terminated store 7951 at 12:06 PM. Operations prepared CDF and D0 for accesses with CDF remaining in access through to the end of the shift. RF technicians replaced the modulator for MIRF1. The readback for the I- Source pressure disappeared; operators accessed to attempt a repair, but it was not successful. The Recycler vacuum valves in the 800 line closed when an argon gas instability caused the ion pumps to trip off. The resulting vacuum cascade caused the loss of the stash. During the evening shift, CDF remained in access until TeV studies began at 6:16 PM; these studies lasted throughout the shift. Operators fired the Recycler TSPs in the 800 region. Antiproton transfers to the Recycler resumed once vacuum recovered. Operations disabled the T-977 beam at 10:37 PM. Wednesday July 7 The midnight shift began with Operations monitoring the stash (170.5E10), with Pbar stacking (11.1E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations reported that the TeV studies ended and operators put the TeV into shot setup at 3:20 AM. Operations established store 7956 at 4:25 AM with an AIL of 235.3E30. Operations established beam to T-977 after the store was in, but disabled the beam at 7:26 AM for alignment work. The Plans for Wednesday The plans for today are to stack, stash, and store. |
|
More Information Comments and Suggestions |
| last modified 07/08/2010 email Fermilab |
| Security, Privacy, Legal | |